# SemSteDiff: Generative Diffusion Model-based Coverless Semantic Steganography Communication

### v1

@@ Abstract
Semantic communication (SemCom), as a novel paradigm for future communication systems, has recently attracted much attention due to its superiority in communication efficiency. However, similar to traditional communication, it also suffers from eavesdropping threats. Intelligent eavesdroppers could launch advanced semantic analysis techniques to infer secret semantic information. Therefore, some researchers have designed Semantic Steganography Communication (SemSteCom) scheme to confuse semantic eavesdroppers. However, the state-of-the-art SemSteCom schemes for image transmission rely on the pre-selected cover image, which limits the universality. To address this issue, we propose a Generative Diffusion Model-based Coverless Semantic Steganography Communication (SemSteDiff) scheme to hide secret images into generated stego images. The semantic related private and public keys enable legitimate receiver to decode secret images correctly while the eavesdropper without completely true key-pairs fail to obtain them. Simulation results demonstrate the effectiveness of the plug-and-play design in different Joint Source-Channel Coding (JSCC) frameworks. The comparison results under different eavesdroppers’ threats show that, when Signal-to-Noise Ratio (SNR) = 0 dB, the peak signal-to-noise ratio (PSNR) of the legitimate receiver is 4.14 dB higher than that of the eavesdropper.
Index Terms:
Semantic communications, image steganography, generative diffusion model, eavesdropping.
@@ I-A Background
With the deep convergence of Artificial intelligence (AI) and communication technologies, the sixth-generation mobile communication system (6G) is expected to connect people, machines, things, and intelligence together to achieve “Internet of Intelligence”. As a pivotal direction for 6G, intellicise (intelligent and concise) wireless network, supported by semantic communication (SemCom), can enable more intelligent, efficient, and low-complexity system [1]. In contrast to traditional communication, SemCom aims to transmit the meaning of messages, called semantic features, rather than transmitting specific bits, thus significantly improving communication efficiency and reducing bandwidth requirements [2, 3]. However, due to the openness of wireless channels, SemCom is vulnerable to eavesdropping threats as traditional communication [4]. Signals transmitted over channel can be intercepted by unauthorized receivers operating within the same frequency band. Once eavesdroppers access to semantic decoders, they could further obtain the private information even under poor channel conditions [5]. Therefore, how to defend against semantic eavesdroppers has become a significant issue.
@@ I-B Secure SemCom
Inspired by convert communication, researchers have recently studied Semantic Steganography Communication (SemSteCom) by introducing steganography technology into SemCom [16, 17, 18, 19, 20, 21, 22]. By embedding the secret information into cover information, steganography provides a new defense method for SemCom to achieve the target of “invisible encryption”. Long et al.[16] introduced a Linguistic steganography scheme, which enhances text imperceptibility and semantic coherence using a knowledge graph to guide secret encoding. Li et al.[17] proposed a multi-modal steganography scheme that hides texts into images to achieve secure SemCom. Towards image transmission, Tang et al.[18] proposed an Invertible Neural Network (INN)-based SemStemCom scheme, which covertly embeds the semantic representation of a private image into the channel input of a host image, yet the eavesdropper can only detect the host image. Besides, Huo et al.[20] constructed a Generating Adversarial Networks (GAN)-based SemSteCom scheme, considering both pixel-level and semantic-level distortions to enhance SemCom security. Also, Chen et al.[21] embedded protected information into the transmission process through a coding-guided jamming strategy. It adapts a two-layer superposition coding structure to prevent eavesdroppers from extracting the privacy source contents.
@@ I-C Motivations and Contributions
Although state-of-the-art SemSteCom schemes embed secret images into cover images to deceive eavesdroppers, they still suffer from the following limitations. First, the hide of secret images is constrained by the size and quality of the cover images. This implies that transmitting a high-dimensional secret image requires a cover image with sufficiently high resolution and semantic characteristics[23]. Second, the cover-edited method faces the risk of robustness. If the cover image suffers from channel noise, compression, and non-linear transformations, it could damage the information of restored secret image[24]. Moreover, the embedding of hidden semantic features could cause statistical inconsistency. Once eavesdroppers employ advanced steganalysis techniques, such potential anomalies may serve as critical cues for detecting semantic steganography [25].
@@
Figure 1: Comparison between existing SemSteCom schemes and the proposed SemSteDiff scheme. (a) introduces the existing cover-edited SemSteCom schemes, which depends on embedding secret messages into cover images to obtain stego images. (b) introduces proposed coverless SemSteDiff scheme, which does not rely on cover images and generates stego images directly.
@@ II-A Network Model
  • Key Agreement Center: The key agreement is responsible for private and public key storage and distribution. Different from traditional public key infrastructure (PKI), where the key pair is almost fixed numbers, our key agreement center updates key pair based on the transmission information, which is task-driven and content-dependent. The private key is determined by the content of secret image while private key is modified from private key. It only sends correct key-pairs to legitimate users to ensure secure key management.
  • Physical Channel: Additive white Gaussian noise (AWGN) is modeled as the physical channel [27]. The noise component $n$ follows a complex Gaussian distribution $\mathcal{CN}(0,\sigma_{w}^{2})$, where $\sigma_{w}^{2}$ indicates the noise power. Hence, the stego images transmitted through channel as follow:
@@
    where $h$ is the coefficients of the physical channel.
  • Eavesdropper: Eavesdropper attempts to intercept transmitted semantic information. Assumed to have the access to semantic decoder, Eavesdropper aims to recover secret images from steganographic semantic representation.
@@
$\mathbf{x}_{s}$} Scret image
@@
II-B Overview of the Proposed SemSteDiff Scheme
Figure 2: The overview framework of SemSteDiff, where BLIP-based private key extractor obtains textual description of secret images as private keys, LLM-based public key generator produces public keys in pairs with private keys, conditional diffusion model-based coverless steganography module embeds keys’ charateristic into attention mechanism to generate relative stego images, and JSCC-based semantic codec module achieves SemCom.
The proposed SemSteDiff mainly comprises three modules, including the BLIP-based semantic key generation, conditional diffusion-based stego image generation, and JSCC-based semantic codec modules. The main parameters are illustrated in Table II. The steps of SemSteDiff are presented in Algorithm 1, with detailed descriptions as following.
@@ II-B1 BLIP-based Semantic Key Generation Module
where $K_{\text{priv}}=(K_{1},K_{2},\dots,K_{T})$ is a sequence of semantic tokens. To enable secure coverless steganography, private key $K_{\text{pub}}$ is modified using a LLM to obtain a natural-language-style public key as
@@ II-B2 Conditional Diffusion Model-based Coverless Steganography Module
Then, the VAE-based encoder converts latent domain $\hat{\mathbf{z}}_{0}$ into bit domain to reconstruct stego image $\mathbf{x}_{\text{stego}}$.
@@
At the receiver, we employ the same steps as transmitter while exchanging the order of $K_{\text{priv}}$ and $K_{\text{pub}}$ to achieve the reverse process by
@@ III-A BLIP for Private Key Generation
Motivated by [24] and [31], we employ BLIP model as a private key generator to extract semantic features from images and use them as conditional inputs to guide the diffusion model in generating semantically consistent stego images. Compared to randomly generated dgital keys, the natural language description itself as the key has obvious advantages. First, such keys are the actual content descriptions of images, enabling precise guidance of diffusion to stego images, which has strong stability. Second, since each image corresponds to a different and content-specific description, the generated keys are unique and random, thereby enhancing the security of the system. Additionally, BLIP specializes in matching image-text pairs and expressing complex scenes. It can avoid subjective biases in manual design while preventing the risk of key leakage caused by human errors.
@@
As illustrated in Alogrithm 2, the private key generation scheme encapsulates the semantic representation of secret images and functions as a cross-modal embedding for secure SemCom. The semantic key $K_{\text{priv}}$ is adaptively derived from the content of each individual image. This specific encoding ensures that private keys remain semantically aligned with the various visual inputs.
@@ III-B LLM for Public Key Generation
The purpose of a coverless steganography task is to generate stego images that do not arouse suspicion of eavesdroppers. Under a specific communication background, eavesdroppers may have a general expectation of the transmitted contents, which leads to a certain restriction of public keys. Due to this situation, the public key must be semantically related to the private key but with different specific content, ensuring the generative diffusion scene guided by public key has similar distributions with secret image, yet remaining content-independent. Even if an eavesdropper knows the class of the transmitted content, the generated stego image will not reveal the true secret information. Based on high requirements for the public keys, we introduce LLMs to generate public key prompts due to their stronger contextual understanding, linguistic diversity, and control capability[33]. In the previous work [34], the LLM is used for converting images to a textual descriptions, then the texts are encrypted and transmitted to the receiver, and finally the receiver decrypts the texts and recovers the original images based on the content. In contrast, we serve the LLM-generated prompts as public keys directly, without additional encryption. This semantics-as-key design simplifies the communication process while still preserving the security of the hidden information.
@@
By utilizing LLM’s generation diversity, our scheme can finely tune private key prompts to get public key shown in Alogrithm 3, leading to a generative diffuison image that has a similar distributions with secret images but remain content-independent, which can further enhance the semantic covertness of the stego image.
@@ IV Conditional Diffusion Model-based Coverless Steganography Module
Figure 4: Process of conditional diffusion model-based coverless steganography. First, the secret image enocodes from pixel space into latent space by VAE encoder. Second, the secret latent vector adds noise under the guidance of private key using attention mechanism. Then, the public key guides the noisy vector to generate the stego latent vector. Finally, the VAE decoder deocde the latent vector into stego image.
@@ IV-A VAE-based Latent Steganography Framework
where $\epsilon$ is a standard Gaussian noise vector. The sampled latent variable $\mathbf{z}_{s}$ serves as the starting point for the following diffusion process. After steganography and semantic transmisson, the variational decoder $\mathcal{D}_{\text{VAE}}(\cdot)$ transforms the latent variable $\mathbf{z}^{\prime}_{s}$ back into the bit space to reconstruct the secret image by
@@
As illustrated in Alogrithm 4, we first employ a VAE[36] to encode an input image $\mathbf{x}$ into a latent representation $\mathbf{z}$ that captures its semantic feature. This latent representation not only retains the semantic core of the image but also suppresses redundant pixel-level details, offering a more compact and secure encoding. Moreover, the latent variable provides a stable and controllable foundation for the subsequent key-guided diffusion generation.
@@ IV-B Key-Guided Stego Image Generation
To effectively integrate the key prompt into the diffusion process, we do not simply concatenate it with image features but instead using an attention mechanism for semanitc alignment. The attention mechanism aims to determine “where to pay attention” and “what to pay attention”, thereby establishing a precise correspondence between image contents and key conditions[37].
@@
The generated keys are embedded into latent space by Contrastive Language–Image Pre-training (CLIP) model as (27), which is a multimodal neural network to extract textual embedding [38], given as follows.
@@
The embeded key prompt $E_{c}$ is projected to obtain keys $K$ and values $V$ and the sampled latent variable $\mathbf{z}_{t}$ is used to calculate the query $Q$ by
@@
where $W_{Q}$, $W_{K}$ and $W_{V}$ are trainable parameters, $t$ is the sampling step [37, 39]. The semantic feature correlation is computed by
@@
In addition, we adopt classifier-free guidance[41] to enhance the flexibility of semantic control. The attention output is fused into the latent feature and passed through a noise prediction model $f\theta$. The model predicts conditional noise $\hat{\varepsilon}_{\text{cond}}$ and unconditional noise $\hat{\varepsilon}_{\text{uncond}}$ by (30) and (31) separately.
@@
where $E_{\emptyset}$ is the null embedding vector. Then, conditional and unconditional noise are weighted using a guidance scale factor $\beta$ to balance semantic consistency and image realism:
@@
Unlike traditional explicit embedding steganography methods, our approach does not modify existing images but instead of performing steganography directly during image generation. As illustrated in Algorithm 5, by conditional sampling process based on semantic keys, the secret information is hidden into the generative path while the public information controlls the content of stego images. This approach not only offers strong security and concealment, but also provides deep support for key recovery and verification at the semantic level.

### v2

@@ Abstract
Semantic communication (SemCom), as a novel paradigm for future communication systems, has recently attracted much attention due to its superiority in communication efficiency. However, similar to traditional communication, it also suffers from eavesdropping threats. Intelligent eavesdroppers could launch advanced semantic analysis techniques to infer secret semantic information. Therefore, some researchers have designed Semantic Steganography Communication (SemSteCom) schemes to confuse semantic eavesdroppers. However, the state-of-the-art SemSteCom schemes for image transmission rely on the pre-selected cover image, which limits the generalization. To address this issue, we propose a Generative Diffusion Model-based Coverless Semantic Steganography Communication (SemSteDiff) scheme to hide secret images into generated stego images. The semantic related private and public keys enable legitimate receiver to decode secret images correctly while the eavesdropper without the completely correct key-pairs fail to obtain them. Simulation results demonstrate the effectiveness of the plug-and-play design in different Joint Source-Channel Coding (JSCC) frameworks. Results under different eavesdropping settings show that, when Signal-to-Noise Ratio (SNR) = 0 dB, the peak signal-to-noise ratio (PSNR) of the legitimate receiver is 4.14 dB higher than that of the eavesdropper.
@@ I-A Background
With the deep convergence of Artificial intelligence (AI) and communication technologies, the sixth-generation mobile communication system (6G) is expected to connect people, machines, things, and intelligence together to achieve “Internet of Intelligence”. As a pivotal direction for 6G, intellicise (intelligent and concise) wireless network, supported by semantic communication (SemCom), can enable a more intelligent, efficient, and low-complexity system [30]. In contrast to traditional communication, SemCom aims to transmit the meaning of messages, called semantic features, rather than transmitting specific bits, thus significantly improving communication efficiency and reducing bandwidth requirements [35, 6].
Numerous researchers have devoted their efforts to SemCom and proposed various key technologies to facilitate its application in 6G, such as generative AI-aided semantic successive refinement [52], semantic-forward relaying [21], and adaptive NOMA-based interference-aware transmission [48]. For instance, Lin et al. [22] investigated semantic interference cancellation for wireless networks, demonstrating that semantic-level signal processing can effectively mitigate co-channel interference and enhance transmission reliability in multi-user 6G scenarios. While SemCom is vulnerable to eavesdropping threats due to the openness of wireless channels [36], making security becomes a significant topic. Signals transmitted over channel can be intercepted by unauthorized receivers operating within the same frequency band. Once eavesdroppers gain access to semantic decoders, they could further obtain the private information even under poor channel conditions [28]. Therefore, how to defend against semantic eavesdroppers has become a significant issue.
@@ I-B Secure SemCom
Inspired by covert communication, researchers have recently studied Semantic Steganography Communication (SemSteCom) by introducing steganography technology into SemCom [24, 20, 39, 40, 14, 5, 29]. By embedding the secret information into cover information, steganography provides a new defense method for SemCom to achieve the target of “invisible encryption”. Long et al.[24] introduced a Linguistic steganography scheme, which enhances text imperceptibility and semantic coherence using a knowledge graph to guide secret encoding. Li et al.[20] proposed a multi-modal steganography scheme that hides texts into images to achieve secure SemCom. Towards image transmission, Tang et al.[39] proposed an Invertible Neural Network (INN)-based SemSteCom scheme, which covertly embeds the semantic representation of a private image into the channel input of a host image, yet the eavesdropper can only detect the host image. Besides, Huo et al.[14] constructed a Generative Adversarial Networks (GAN)-based SemSteCom scheme, considering both pixel-level and semantic-level distortions to enhance SemCom security. Also, Chen et al.[5] embedded protected information into the transmission process through a coding-guided jamming strategy. It adapts a two-layer superposition coding structure to prevent eavesdroppers from extracting the privacy source contents.
@@ I-C Motivations and Contributions
Although state-of-the-art SemSteCom schemes embed secret images into cover images to deceive eavesdroppers, they still suffer from the following limitations. First, the hiding of secret images is constrained by the size and quality of the cover images. This implies that transmitting a high-dimensional secret image requires a cover image with sufficiently high resolution and semantic characteristics[55]. Second, the cover-edited method faces the robustness risk. If the cover image suffers from channel noise, compression, and non-linear transformations, it could damage the information of restored secret image[51]. Moreover, the embedding of hidden semantic features could cause statistical inconsistency. Once eavesdroppers employ advanced steganalysis techniques, such potential anomalies may serve as critical cues for detecting semantic steganography [37].
@@
Figure 1: Comparison between existing SemSteCom schemes and the proposed SemSteDiff scheme. (a) introduces the existing cover-edited SemSteCom schemes, which depend on embedding secret messages into cover images to obtain stego images. (b) introduces proposed coverless SemSteDiff scheme, which does not rely on cover images and generates stego images directly.
@@ II-A Network Model
  • Key Agreement Center: The key agreement center is responsible for private and public key storage and distribution. Different from traditional public key infrastructure (PKI), where the key pairs are almost fixed numbers, our key agreement center updates key pairs based on the transmission information, which is task-driven and content-dependent. The private key is determined by the content of secret image, while the public key is generated from the private key by modifying the core semantic meaning. The interaction between legitimate users and the key management center is assumed to be conducted through secure channels [43].
  • Physical Channel: Additive white Gaussian noise (AWGN) is modeled as the physical channel [45]. The noise component $n$ follows a complex Gaussian distribution $\mathcal{CN}(0,\sigma_{w}^{2})$, where $\sigma_{w}^{2}$ indicates the noise power. Hence, the stego images transmitted through channel as follows:
@@
    where $h$ is the coefficient of the physical channel.
  • Eavesdropper: Eavesdropper attempts to intercept transmitted semantic feature $\mathcal{S}’_{Eve}$ from channel by
    $\mathcal{S}’_{Eve}=h*\mathcal{S}+n,$ (2)
    where $\mathcal{S}$ is the original semantic feature. Assumed to have access to semantic decoder $\mathcal{D}_{sem}(\cdot)$, eavesdropper tries to recover the transmitted image $\mathbf{x}’_{trans}$ by
    $\mathbf{x}’_{trans}=\mathcal{D}_{sem}(\mathcal{S}^{\prime}_{Eve}).$ (3)
    Based on the covertness of SemSteCom, the eavesdropper typically assumes that the decoded stego image $\mathbf{x}’_{trans}$ corresponds to the real secret image. To further assess the security of the proposed framework, we also consider a worst-case scenario that the eavesdropper is aware that steganography is employed and actively attempts to break the stego image by LLM or other advanced generative models, given by
[…]
@@
$\mathbf{x}_{s}$} Secret image
@@
II-B Overview of the SemSteDiff Scheme
Figure 2: The overview framework of SemSteDiff, where BLIP-based private key extractor obtains textual description of secret images as private keys, LLM-based public key generator produces public keys in pairs with private keys, conditional diffusion model-based coverless steganography module embeds keys’ characteristic into attention mechanism to generate relative stego images, and JSCC-based semantic codec module achieves SemCom.
The proposed SemSteDiff mainly comprises three modules, including the BLIP-based semantic key generation, conditional diffusion-based stego image generation, and JSCC-based semantic codec modules. The main parameters are illustrated in Table II. The steps of SemSteDiff are presented in Algorithm 1, with detailed descriptions as follows.
@@ II-B1 BLIP-based Semantic Key Generation Module
where $K_{\text{priv}}=(K_{1},K_{2},\dots,K_{T})$ is a sequence of semantic tokens. To enable secure coverless steganography, private key $K_{\text{priv}}$ is modified using a LLM to obtain a natural-language-style public key as
@@ II-B2 Conditional Diffusion Model-based Coverless Steganography Module
Then, the VAE-based decoder converts latent domain $\hat{\mathbf{z}}_{0}$ into bit domain to reconstruct stego image $\mathbf{x}_{\text{stego}}$.
@@
At the receiver, we employ the same steps as the transmitter while exchanging the order of $K_{\text{priv}}$ and $K_{\text{pub}}$ to achieve the reverse process by
@@ III-A BLIP for Private Key Generation
Motivated by [51] and [50], we employ BLIP model as a private key generator to extract semantic features from images and use them as conditional inputs to guide the diffusion model in generating semantically consistent stego images. Compared to randomly generated digital keys, the natural language description itself as the key has obvious advantages. First, such keys are the actual content descriptions of images, enabling precise guidance of diffusion to stego images, which has strong stability. Second, since each image corresponds to a different and content-specific description, the generated keys are unique and random, thereby enhancing the security of the system. Additionally, BLIP specializes in matching image-text pairs and expressing complex scenes. It can avoid subjective biases in manual design while preventing the risk of key leakage caused by human errors.
@@
As illustrated in Algorithm 2, the private key generation scheme encapsulates the semantic representation of secret images and functions as a cross-modal embedding for secure SemCom. The semantic key $K_{\text{priv}}$ is adaptively derived from the content of each individual image. This specific encoding ensures that private keys remain semantically aligned with the various visual inputs.
@@ III-B LLM for Public Key Generation
The purpose of a coverless steganography task is to generate stego images that do not arouse suspicion of eavesdroppers. Under a specific communication background, eavesdroppers may have a general expectation of the transmitted contents, which leads to a certain restriction of public keys. Due to this situation, the public key must be semantically related to the private key but with different specific content, ensuring the generative diffusion scene guided by public key has similar distributions with secret image, yet remaining content-independent. Even if an eavesdropper knows the class of the transmitted content, the generated stego image will not reveal the true secret information. Based on high requirements for the public keys, we introduce LLMs to generate public key prompts due to their stronger contextual understanding, linguistic diversity, and control capability[31]. In the previous work [3], the LLM is used for converting images to a textual descriptions, then the texts are encrypted and transmitted to the receiver, and finally the receiver decrypts the texts and recovers the original images based on the content. In contrast, we serve the LLM-generated prompts as public keys directly, without additional encryption. Specific LLM-required prompt is designed to control the generation process only preserving general category and background of private key sentence, while detailed descriptions related to objects, attributes, or behavior are intentionally replaced with irrelevant counterparts, which ensures the public key remains semantic relative to private key without any leakage. This semantics-as-key design simplifies the communication process while still preserving the security of the hidden information.
@@
To ensure that the public key remains semantically related to the private key while remaining sufficiently distinct to mislead eavesdroppers, several public keys towards on one private key are generated to avoid generation collapsing. The semantic similarity between public key-guided stego images and secret images, as well as the correlation between public keys and stego images are both considered by evaluating embedding-based similarity such as Contrastive Language–Image Pre-training (CLIP) score[32] to comprehensively to choose the optimal one. It jointly prevent semantic leakage caused by over-specific descriptions and avoid uninformative public keys, which further reduces the likelihood of generating similar public keys across different secret images, improving the robustness and security of the LLM-based key generation process in SemSteDiff.
@@
By utilizing LLM’s generation diversity, our scheme can finely tune private key prompts to get public key shown in Algorithm 3, leading to a generative diffusion image that has a similar distributions with secret images but remain content-independent, which can further enhance the semantic covertness of the stego image.
@@ IV Conditional Diffusion Model-based Coverless Steganography Module
Figure 4: Process of conditional diffusion model-based coverless steganography. First, the secret image encodes from pixel space into latent space by VAE encoder. Second, the secret latent vector adds noise under the guidance of private key using attention mechanism. Then, the public key guides the noisy vector to generate the stego latent vector. Finally, the VAE decoder decodes the latent vector into stego image.
@@ IV-A VAE-based Latent Steganography Framework
where $\epsilon$ is a standard Gaussian noise vector. The sampled latent variable $\mathbf{z}_{s}$ serves as the starting point for the following diffusion process. After steganography and semantic transmission, the variational decoder $\mathcal{D}_{\text{VAE}}(\cdot)$ transforms the latent variable $\mathbf{z}^{\prime}_{s}$ back into the bit space to reconstruct the secret image by
@@
As illustrated in Algorithm 4, we employ a VAE[18] to encode an input image $\mathbf{x}$ into a latent representation $\mathbf{z}$ that captures its semantic feature. The steganographic capacity is constrained by the VAE-based latent space $\mathcal{Z}\in\mathbb{R}^{h\times w\times c}$, where each secret image is mapped into a compact latent vector. The capacity per generated stego image is simply calculated as $C_{L}=(h\times w\times c)\times b$ bits, where $b$ represents the numerical precision. This latent representation retains the semantic core of the image and suppresses redundant pixel-level details to offer a more compact encoding, which also provides a stable and controllable foundation for the subsequent key-guided diffusion generation.
@@ IV-B Key-Guided Stego Image Generation
To effectively integrate the key prompt into the diffusion process, we do not simply concatenate it with image features but instead using an attention mechanism for semantic alignment. The attention mechanism aims to determine “where to pay attention” and “what to pay attention”, thereby establishing a precise correspondence between image contents and key conditions[42].
@@
The generated keys are embedded into latent space by CLIP model, which is a multimodal neural network to extract textual embedding [32],given as follows.
@@
The embedding key prompt $E_{c}$ is projected to obtain keys $K$ and values $V$ and the sampled latent variable $\mathbf{z}_{t}$ is used to calculate query $Q$ by
@@
where $W_{Q}$, $W_{K}$ and $W_{V}$ are trainable parameters, $t$ is the sampling step [42, 15]. It combines semantic keys with diffusion process by controling every sampling steps. The semantic feature correlation is computed by
@@
In addition, we adopt classifier-free guidance[12] to enhance the flexibility of semantic control. The attention output is fused into the latent feature and passed through a noise prediction model $f_{\theta}$. The model predicts conditional noise $\hat{\varepsilon}_{\text{cond}}$ and unconditional noise $\hat{\varepsilon}_{\text{uncond}}$ by (33) and (34) separately.
@@
where $E_{\emptyset}$ is the null embedding vector, and Key is applied to compute the $E_{c}$ in (30). Then, conditional and unconditional noise are weighted using a guidance scale factor $\beta$ to balance semantic consistency and image realism:
@@
Unlike traditional explicit embedding steganography methods, our approach does not modify existing images but instead of performing steganography directly during image generation. As illustrated in Algorithm 5, by conditional sampling process based on semantic keys, the secret information is hidden into the generative path while the public information controls the content of stego images.